OpenStudy (anonymous):

Solve (2b + 1)^2 = 25

OpenStudy (callisto):

Take square root of both sides 2b + 1 = 5 or 2b + 1 = -5 Now, solve the 2 equations, can you?

OpenStudy (shubhamsrg):

either you can expand LHS and form a quadratic equation and solve,,or you can see that 25= 5^2 or (-5)^2 ..then you can take root of both sides and solve for b,,you should get 2 values..

OpenStudy (anonymous):

i need the values

OpenStudy (callisto):

Have you tried?

OpenStudy (anonymous):

yea

OpenStudy (shubhamsrg):

-3 and 2 ..you can check by hit and trial also..this makes it an oral question.

OpenStudy (anonymous):

so just -3 and 2 or x=-3 x=2 im askin cause idk

OpenStudy (shubhamsrg):

ydk !!?? hmmn,,anyways its b=-3 and b=2..

OpenStudy (anonymous):

thanks thats all I wanted

OpenStudy (apoorvk):

@shubhamsrg unfortunately, hit-and-trial though handy, doesn't always work :) Simplify the equation, bring every term to the same side, and then factorize the quadratic you get.

OpenStudy (shubhamsrg):

well yes it dosent work every time,,but see the given equation,,saves time with hit and trial..umm..